OpenStudy (anonymous):

find the volume of a regular tetrahedron whose altitude measures 10cm.

OpenStudy (anonymous):

what formula did you use?

OpenStudy (anonymous):

hmm thank you for the answer

OpenStudy (wolf1728):

Going by what wikipedia says: http://en.wikipedia.org/wiki/Tetrahedron the height of the tetrahedron (altitude?) is sq root (.66666666) * edge length edge length = height / sq root (2/3) edge length = 10 / 0.8164965809 edge length = 12.24745 Volume = (edge length)^3 /(6 * sq root (2) ) Volume = 1,837.12 / 8.485 Volume = 216.51 cubic cm

OpenStudy (anonymous):

wolf1728 can u explain further?

OpenStudy (wolf1728):

Notice I changed the calculations slightly Also the formula for volume is (edge length) ^ 3 / (6 * square root (2))

OpenStudy (anonymous):

tnx

OpenStudy (wolf1728):

So, everything looks okay now?

OpenStudy (anonymous):

im a bit confused but ill just understand it

OpenStudy (wolf1728):

You need the edge length to calculate the volume. The edge length equals the altitude (or height) divided by the square root of (2/3) The volume = (edge^3) / (6 * square root (2))

OpenStudy (anonymous):

okay, tnx for the help :)

OpenStudy (wolf1728):

okay - glad to help out